OpenStudy (anonymous):

You have a CD with 15 songs, and you want to play 3 of them for your friends. How many different arrangements of 3 songs are possible? (Count two arrangements as different if they have the same 3 songs in a different order.) A)455 B)45 C)2,730 D)3,375

OpenStudy (cwrw238):

i think this is 15!/3!

OpenStudy (anonymous):

15 songs for the first, 14 for the the second 13 for the last (of the three) = 15*14*13 = 2,730

OpenStudy (cwrw238):

yup - my stats aint brilliant! micha is right

OpenStudy (cwrw238):

if the order of the three choices didnt matter it would be 2730/6 = 455

OpenStudy (anonymous):

thank you!